# A quantitative assessment of Geant4 for predicting the yield and distribution of positron-emitting fragments in ion beam therapy

### v1

@@ Abstract
Materials and Methods: Phantoms of polyethylene, gelatin or poly(methyl methacrylate) were irradiated with monoenergetic carbon and oxygen ion beams. Post-irradiation, 4D PET images were acquired and parent ${}^{11}$, ${}^{10}$ and ${}^{15}$ radionuclides contributions in each voxel were determined from the extracted time activity curves. Next, the experimental configurations were simulated in Geant4 Monte Carlo versions 10.0 to 11.1, with three different fragmentation models - binary ion cascade (BIC), quantum molecular dynamics (QMD) and the Liege intranuclear cascade (INCL++) - 30 model-version combinations. Total positron annihilation and parent isotope production yields predicted by each simulation were compared between simulations and experiments using normalised mean squared error and Pearson cross-correlation coefficient. Finally, we compared the depth of maximum positron annihilation yield and the distal point at which positron yield decreases to 50% of peak between each model and the experimental results.
@@ 1 Introduction
In addition to the normalised mean squared error (NMSE) metric used in the previous study, three additional metrics - the Pearson cross-correlation coefficient (CCC), the depth of the positron annihilation peak, and the depth at which positron annihilation intensity has decreased to 50% of peak are also used to compare the shape of the predicted positron-emitting fragment distribution to the experimentally-measured distribution. The experimental absolute yields of the dominant positron-emitting fragment species (${}^{10}$, ${}^{11}$ and ${}^{15}$) created during irradiation of gelatin, PMMA and polyethylene block phantoms with beams of ${}^{12}$ ions with energies of 148.5, 290.5 and 350 MeV/u and ${}^{16}$ ions with energies of 148 and 290 MeV/u were obtained across the full width at tenth maximum (FWTM) of the beam from a previous study conducted using the OpenPET scanner at HIMAC, Japan [12, 30]. Akamatsu2019 results in each of the three target materials and 5 ion/energy combinations were then compared to those predicted in equivalent simulations performed in Geant4 using each of hadronic fragmentation models (BIC, QMD and INCL++) across the ten evaluated Geant4 versions for a total of 150 unique target/ion/energy/version/model test conditions. The total positron yields and yields of the individual positron-emitting fragment species from each model and Geant4 version were then compared to experimental measurements using each of the chosen metrics at the entrance, build-up and Bragg peak, and tail regions.
@@ 2 Materials and Methods
The experimental methods used to estimate the total positron annihilation profile and activity of the dominant positron-emitting fragment isotopes (${}^{11}$, ${}^{10}$ and ${}^{15}$) are briefly summarised in Section 2.1. Equivalent simulation configurations were constructed for each Geant4 version under test, and the total positron annihilation profile and activity of ${}^{11}$, ${}^{10}$ and ${}^{15}$ were predicted for each beam ion/energy, target material, hadronic inelastic fragmentation model and Geant4 version; the design and parameters of these simulations are described in detail in Section 2.2.
@@ 2.1 Experimental configuration
The experimental data obtained in our 2019 paper were used as the ground truth for this simulation study; a detailed description of the experimental procedures is presented in that paper [30]. In summary, phantoms constructed from either pure PMMA, polyethylene or gelatin (encased in a thin-walled PMMA container), each with dimensions of 100 mm×$\times$100 mm×$\times$300 mm, were irradiated with monoenergetic carbon or oxygen ion beams of various energies - three for carbon ions and two for oxygen (see Table 1). Positron annihilation profiles (with respect to depth in the target) were obtained using the whole-body DOI-PET scanner prototype developed at QST [12]. These profiles were decomposed into the individual population of each of the dominant parent positron-emitting fragments (${}^{11}$, ${}^{10}$ and ${}^{15}$) at $t=0$ (end of irradiation period) by fitting the observed time-decay curves in each voxel to a multiexponential decay model.
@@
Ion Energy (MeV/u)
$\sigma_{x}$ (mm)
$\sigma_{y}$ (mm)
Beam flux (pps)
${}^{12}$
148.5
2.77
2.67
@@ 2.2 Simulation parameters
[…]
@@ 2.3 Evaluation methods
For each version of Geant4, phantom, beam type and energy, the NMSE and CCC were calculated for both total annihilation photon yield profiles and also for the profiles of the three main positron-emitting fragment species (${}^{10}$, ${}^{11}$ and ${}^{15}$). The calculation was repeated for each of the $N_{reg}$ regions (entrance, build-up and Bragg peak, and tail regions). The NMSE and the CCC were then compared across all evaluated Geant4 versions for each region, phantom material and beam type.
A total of 5 energy/ion combinations are evaluated (carbon ions at three energies and oxygen ions at two energies). For oxygen ions, three target materials (gelatine, PMMA and polyethylene) are evaluated for total positron annihilation yield and ${}^{11}$/${}^{10}$/${}^{15}$ yield. For carbon ions, the same three target materials are evaluated for total positron annihilation yield and ${}^{11}$/${}^{10}$ yield and two for ${}^{15}$ yield (polyethylene is omitted since it is not possible to produce ${}^{15}O$ fragments with a ${}^{12}C$ ion beam and a PE target which only contains carbon and hydrogen). Thus, there are a total of 15 cases evaluated for total positron annihilation, ${}^{11}$ yield and ${}^{10}$ yield, while there are 12 cases evaluated for ${}^{15}$ yield.
@@ 3.1 Entrance region
10
6
0
0
11
3
2
0
0
2
6
2
0
10.1
6
0
0
11
3
2
0
0
3
6
2
0
10.2
5
0
0
11
3
2
0
0
3
5
2
0
10.3
6
0
0
6
0
0
5
3
6
4
2
0
10.4
6
0
0
1
0
0
2
3
6
9
2
0
10.5
2
0
0
0
0
0
2
5
8
3
1
0
10.6
4
0
0
1
0
0
0
0
5
5
4
0
10.7
3
0
0
1
0
0
0
0
5
5
2
0
11
4
0
0
1
0
0
0
0
5
5
2
0
11.1
4
0
0
0
0
0
0
0
5
5
2
0
For the entrance region, the BIC model implemented in Geant4 versions 10, 10.1, 10.3 and 10.4 provided the (equal) lowest NMSE of the yields of total positron annihilation in 5 out of 15 cases. The BIC model in Geant4 10, 10.1 and 10.2 also provided the (equal) lowest NMSE for ${}^{11}$ fragment production (11/15 cases), whereas for ${}^{10}$ the best version/model combination was 10.5/INCL (8/15 cases) and for ${}^{15}$ it was 10.6/BIC (9/12 cases).
@@
Results for individual radionuclides were also mixed, with 10/BIC, 10.1/BIC, 10.4/BIC and 10.5/INCL achieving the threshold in 4/15 cases for ${}^{11}$, 10-10.4/BIC and all versions with QMD reaching the threshold in 2/15 cases for ${}^{10}$, and all versions with BIC and 10/INCL, 10.1/INCL, 10.2/INCL, and 10.5/INCL reaching the threshold for ${}^{15}$.
@@ 3.2 Build-up and Bragg peak region
10
4
0
0
11
6
3
0
0
2
3
2
0
10.1
5
0
0
11
6
3
0
0
2
3
2
0
10.2
11
1
0
14
6
3
0
0
1
5
1
0
10.3
1
0
0
5
0
0
4
1
2
3
0
0
10.4
1
0
0
1
0
0
3
2
2
4
0
0
10.5
0
0
0
0
0
0
3
4
7
1
2
0
10.6
1
0
0
0
0
0
3
0
9
2
1
2
10.7
1
0
0
0
0
0
3
2
7
0
0
1
11
2
0
0
0
0
0
3
2
9
2
0
2
11.1
1
0
0
0
0
0
3
2
9
2
0
2
In the build-up and Bragg peak region, according to the NMSE metric, total positron yield is most accurately predicted by the BIC model in Geant4 version 10.2, being (equal) best in 11/15 cases. This is much higher than the next-best combinations (10.1/BIC with 5/11 cases followed by 10/BIC with 4/11). Similar results are observed for ${}^{11}$ yield, with 10.2/BIC achieving (equal) best performance in 14/15 cases, and 10/BIC and 10.1/BIC each achieving (equal) best results in 11/15 cases; QMD also performs reasonably well in this case with 10, 10.1 and 10.2 achieving wins in 6/15 cases. For ${}^{10}$, 10.6/INCL, 11/INCL and 11.1/INCL are the best performers (each winning in 9/15 cases). Finally, for ${}^{15}$, 10.2/BIC is the best-performing model with 5/12 wins, followed by 10.4/BIC with 4 wins.
@@
10
9
4
3
3
3
2
4
3
6
3
2
3
10.1
9
4
4
3
3
3
4
3
6
3
2
3
10.2
10
8
4
6
6
3
5
4
5
3
3
3
10.3
9
6
6
6
5
3
2
2
2
4
3
3
10.4
6
8
7
6
5
3
1
1
3
4
3
3
10.5
8
8
6
6
5
4
2
2
3
3
3
4
10.6
9
9
6
6
7
6
2
1
4
4
3
4
10.7
6
6
4
3
5
2
2
1
4
3
2
3
11
8
8
6
4
5
4
3
2
4
3
3
3
11.1
6
5
6
2
3
4
3
2
4
3
3
3
Using the Pearson CCC metric, the best-performing version/model combinations for overall positron yield are 10.2/BIC (10/15 cases), followed by 10/BIC, 10.1/BIC, 10.3/BIC, 10.6/BIC and 10.6/QMD (9/15 cases). Generally, BIC performed very well, with all Geant4 versions achieving (equal) best performance in at least 6 cases. ${}^{11}$ yield was best predicted by 10.6/QMD (7/15 cases) however many version/model combinations did well here also, with 10.2/BIC, 10.2/QMD, 10.3/BIC, 10.4/BIC, 10.5/BIC, 10.6/BIC and 10.6/INCL all achieving 6/15 wins. ${}^{10}$ yield was best predicted by 10/INCL and 10.1 INCL (6/15 cases), closely followed by 10.2/BIC and 10.2/INCL which won in 5/15 cases. The best-performing version/model combinations for ${}^{15}$ yield were 10.3/BIC, 10.4/BIC, 10.5/INCL, 10.6/BIC and 10.6/INCL with 4/15 wins each, and all other version/model combinations achieving 2 or 3 wins.
@@
$\mu$ (mm)
$\sigma$ (mm)
max (mm)
$\mu$ (mm)
$\sigma$ (mm)
max (mm)
$\mu$ (mm)
$\sigma$ (mm)
max (mm)
10
1
1.85
6
-0.20
1.69
3
1.10
3.82
10.50
10.1
1
1.85
6
-0.20
1.69
3
1
3.91
10.50
10.2
0.60
1.37
3
-0.60
1.58
-3
0.60
3.39
9
10.3
1.30
1.78
6
-0.30
1.41
-3
2.30
4.12
9
10.4
1.60
1.55
6
-0.10
1.44
-3
4
4.45
10.50
10.5
0.69
0.99
3
0.60
2.03
6
2.20
4.24
9
10.6
0.60
1.37
3
-0.10
1.55
3
0.10
2.50
7.50
10.7
1.20
1.72
4.50
0.60
1.95
4.50
0.70
3.40
10.50
11
1.10
1.65
4.50
0.60
1.95
4.50
0.60
3.09
9
11.1
0.30
1.62
3
-0.30
1.62
3
0
2.78
7.50
@@
$\mu$ (mm)
$\sigma$ (mm)
max (mm)
$\mu$ (mm)
$\sigma$ (mm)
max (mm)
$\mu$ (mm)
$\sigma$ (mm)
max (mm)
10
0.70
1.49
3
0.30
1.41
3
-0.20
1.49
-3
10.1
0.70
1.49
3
0.30
1.41
3
-0.20
1.49
-3
10.2
0.30
1.01
1.50
0
1.13
1.50
-0.50
1.22
-3
10.3
0.50
1.09
3
0.20
1.11
1.50
-0.20
1.37
-3
10.4
0.60
1.11
3
0.30
1.01
1.50
0.10
1.20
-3
10.5
0.35
0.90
1.50
0.20
1.11
1.50
-0.50
1.22
-3
10.6
0.40
0.89
1.50
0.20
1.11
1.50
-0.40
1.33
-3
10.7
1
1.46
3
0.70
1.59
3
-0.10
1.65
3
11
1
1.46
3
0.70
1.59
3
0
1.60
3
11.1
0
1.60
-3
-0.10
1.44
-3
-0.60
1.68
-3
@@ 3.3 Tail region
10
3
2
2
5
9
4
2
2
1
4
4
4
10.1
4
2
2
6
9
4
1
1
2
4
4
4
10.2
12
10
2
11
12
4
1
1
3
7
7
4
10.3
1
1
1
1
1
1
2
4
0
3
4
4
10.4
1
1
1
1
1
1
2
3
0
3
4
3
10.5
1
1
1
1
1
1
3
4
2
2
2
2
10.6
1
2
1
1
1
1
5
3
6
4
10
4
10.7
1
1
1
1
1
1
3
3
5
3
4
3
11
1
1
1
1
1
1
4
3
6
4
4
3
11.1
1
1
1
1
1
1
4
3
5
4
4
3
Using the NMSE metric, 10.2/BIC was the best-performing version/model combination for overall positron yield (12/15 cases), with 10.2/QMD being the second-best (10/15). Results were similar for ${}^{11}$ yield, with the best version/model combinations being 10.2/QMD (12/15 cases) and 10.2/BIC (11/15). For ${}^{10}$, the most wins were obtained by 10.6/INCL and 11/INCL (6/15 cases) followed by 10.6/BIC, 10.7/INCL and 11.1/INCL (5/15 cases). Finally, for ${}^{15}$, the best results were obtained with 10.6/QMD (10/12 cases) followed by 10.2/BIC and 10.2/QMD (7/12 cases).
@@
10
12
11
11
10
11
11
4
3
4
8
7
9
10.1
12
11
11
11
11
11
4 4 4
8
7
9
10.2
12
11
11
10
11
11
4 4 4
8
7
7
10.3
11
11
10
11
11
11
2
2
1
8
7
8
10.4
10
10
10
9
11
10
2
2
0
7
7
8
10.5
11
11
11
11
11
11
2
2
3
7
7
8
10.6
11
11
12
11
11
11
2
3
3
7
7
9
10.7
12
11
12
11
11
12
3
3
3
7
7
9
11
12
11
12
11
11
13
3
3
3
8
8
9
11.1
12
11
12
11
11
13
3
3
3
8
8
9
The Pearson CCC results in the tail region were all very similar across Geant4 versions, with only a few wins separating the best and worst-performing version/model combinations in most instances. All version/models exceeded the threshold of 0.95 for a clear majority of cases for total positron yield as well as ${}^{11}$ and ${}^{15}$ production. For total positron annihilation yield, 10/BIC, 10.1/BIC, 10.2/BIC, 10.6/INCL, 10.7/BIC, 10.7/INCL, 11/BIC, 11/INCL, 11.1/BIC and 11.1/INCL all exceeded the target threshold for 12/15 cases. Even the worst-performing version/model combinations still exceeded the threshold in 10/15 cases. For ${}^{11}$ yield, 11/INCL and 11.1/INCL reached the threshold in 13/15 cases (with the worst-performing combination scoring 9/15 wins). Fewer wins were seen with ${}^{10}$; the best results were obtained with 10/BIC, 10/INCL, 10.1/BIC, 10.1/QMD, 10.1/INCL, 10.2/BIC, 10.2/QMD and 10.2/INCL (4/15 cases). Finally, ${}^{15}$ yield was best predicted by 10/INCL, 10.1/INCL, 10.6/INCL, 10.7/INCL, 11/INCL and 11.1/INCL (9/12 cases) - again, in this case, even the worst-performing version/model combinations exceeded the threshold in 7/12 cases.

### v2

@@ Abstract
Materials and Methods: Phantoms of polyethylene, gelatin or poly(methyl methacrylate) were irradiated with monoenergetic carbon and oxygen ion beams. Post-irradiation, 4D PET images were acquired and parent 11C, 10C and 15O radionuclides contributions in each voxel were determined from the extracted time activity curves. Next, the experimental configurations were simulated in Geant4 Monte Carlo versions 10.0 to 11.1, with three different fragmentation models - binary ion cascade (BIC), quantum molecular dynamics (QMD) and the Liege intranuclear cascade (INCL++) - 30 model-version combinations. Total positron annihilation and parent isotope production yields predicted by each simulation were compared between simulations and experiments using normalised mean squared error and Pearson cross-correlation coefficient. Finally, we compared the depth of maximum positron annihilation yield and the distal point at which positron yield decreases to 50% of peak between each model and the experimental results.
@@ 1 Introduction
In addition to the normalised mean squared error (NMSE) metric used in the previous study, three additional metrics - the Pearson cross-correlation coefficient (CCC), the depth of the positron annihilation peak, and the depth at which positron annihilation intensity has decreased to 50% of peak are also used to compare the shape of the predicted positron-emitting fragment distribution to the experimentally-measured distribution. The experimental absolute yields of the dominant positron-emitting fragment species (10C, 11C and 15O) created during irradiation of gelatin, PMMA and polyethylene block phantoms with beams of 12C ions with energies of 148.5, 290.5 and 350 MeV/u and 16O ions with energies of 148 and 290 MeV/u were obtained across the full width at tenth maximum (FWTM) of the beam from a previous study conducted using the OpenPET scanner at HIMAC, Japan [12, 30]. Akamatsu2019 results in each of the three target materials and 5 ion/energy combinations were then compared to those predicted in equivalent simulations performed in Geant4 using each of hadronic fragmentation models (BIC, QMD and INCL++) across the ten evaluated Geant4 versions for a total of 150 unique target/ion/energy/version/model test conditions. The total positron yields and yields of the individual positron-emitting fragment species from each model and Geant4 version were then compared to experimental measurements using each of the chosen metrics at the entrance, build-up and Bragg peak, and tail regions.
@@ 2 Materials and Methods
The experimental methods used to estimate the total positron annihilation profile and activity of the dominant positron-emitting fragment isotopes (11C, 10C and 15O) are briefly summarised in Section 2.1. Equivalent simulation configurations were constructed for each Geant4 version under test, and the total positron annihilation profile and activity of 11C, 10C and 15O were predicted for each beam ion/energy, target material, hadronic inelastic fragmentation model and Geant4 version; the design and parameters of these simulations are described in detail in Section 2.2.
@@ 2.1 Experimental configuration
The experimental data obtained in our 2019 paper were used as the ground truth for this simulation study; a detailed description of the experimental procedures is presented in that paper [30]. In summary, phantoms constructed from either pure PMMA, polyethylene or gelatin (encased in a thin-walled PMMA container), each with dimensions of 100 mm×$\times$100 mm×$\times$300 mm, were irradiated with monoenergetic carbon or oxygen ion beams of various energies - three for carbon ions and two for oxygen (see Table 1). Positron annihilation profiles (with respect to depth in the target) were obtained using the whole-body DOI-PET scanner prototype developed at QST [12]. These profiles were decomposed into the individual population of each of the dominant parent positron-emitting fragments (11C, 10C and 15O) at $t=0$ (end of irradiation period) by fitting the observed time-decay curves in each voxel to a multiexponential decay model.
@@
Ion Energy (MeV/u) $\sigma_{x}$ (mm) $\sigma_{y}$ (mm) Beam flux (pps)
@@ 2.2 Simulation parameters
The location of each positron annihilation, as well as the identity of the parent isotope which decayed to emit each positron (principally 11C, 10C and 15O), were scored with a resolution of 1.5 mm3 to match the voxel dimensions of the experimental OpenPET image reconstruction output. The pristine positron annihilation profiles were convolved with a 2.3 mm FWHM Gaussian filter to simulate the point spread function of the PET system.
@@ 2.3 Evaluation methods
For each version of Geant4, phantom, beam type and energy, the NMSE and CCC were calculated for both total annihilation photon yield profiles and also for the profiles of the three main positron-emitting fragment species (10C, 11C and 15O). The calculation was repeated for each of the $N_{reg}$ regions (entrance, build-up and Bragg peak, and tail regions). The NMSE and the CCC were then compared across all evaluated Geant4 versions for each region, phantom material and beam type.
A total of 5 energy/ion combinations are evaluated (carbon ions at three energies and oxygen ions at two energies). For oxygen ions, three target materials (gelatine, PMMA and polyethylene) are evaluated for total positron annihilation yield and 11C/10C/15O yield. For carbon ions, the same three target materials are evaluated for total positron annihilation yield and 11C/10C yield and two for 15O yield (polyethylene is omitted since it is not possible to produce ${}^{15}O$ fragments with a ${}^{12}C$ ion beam and a PE target which only contains carbon and hydrogen). Thus, there are a total of 15 cases evaluated for total positron annihilation, 11C yield and 10C yield, while there are 12 cases evaluated for 15O yield.
@@ 3.1 Entrance region
10 6 0 0 11 3 2 0 0 2 6 2 0
10.1 6 0 0 11 3 2 0 0 3 6 2 0
10.2 5 0 0 11 3 2 0 0 3 5 2 0
10.3 6 0 0 6 0 0 5 3 6 4 2 0
10.4 6 0 0 1 0 0 2 3 6 9 2 0
10.5 2 0 0 0 0 0 2 5 8 3 1 0
10.6 4 0 0 1 0 0 0 0 5 5 4 0
10.7 3 0 0 1 0 0 0 0 5 5 2 0
11 4 0 0 1 0 0 0 0 5 5 2 0
11.1 4 0 0 0 0 0 0 0 5 5 2 0
For the entrance region, the BIC model implemented in Geant4 versions 10, 10.1, 10.3 and 10.4 provided the (equal) lowest NMSE of the yields of total positron annihilation in 5 out of 15 cases. The BIC model in Geant4 10, 10.1 and 10.2 also provided the (equal) lowest NMSE for 11C fragment production (11/15 cases), whereas for 10C the best version/model combination was 10.5/INCL (8/15 cases) and for 15O it was 10.6/BIC (9/12 cases).
@@
Results for individual radionuclides were also mixed, with 10/BIC, 10.1/BIC, 10.4/BIC and 10.5/INCL achieving the threshold in 4/15 cases for 11C, 10-10.4/BIC and all versions with QMD reaching the threshold in 2/15 cases for 10C, and all versions with BIC and 10/INCL, 10.1/INCL, 10.2/INCL, and 10.5/INCL reaching the threshold for 15O.
@@ 3.2 Build-up and Bragg peak region
10 4 0 0 11 6 3 0 0 2 3 2 0
10.1 5 0 0 11 6 3 0 0 2 3 2 0
10.2 11 1 0 14 6 3 0 0 1 5 1 0
10.3 1 0 0 5 0 0 4 1 2 3 0 0
10.4 1 0 0 1 0 0 3 2 2 4 0 0
10.5 0 0 0 0 0 0 3 4 7 1 2 0
10.6 1 0 0 0 0 0 3 0 9 2 1 2
10.7 1 0 0 0 0 0 3 2 7 0 0 1
11 2 0 0 0 0 0 3 2 9 2 0 2
11.1 1 0 0 0 0 0 3 2 9 2 0 2
In the build-up and Bragg peak region, according to the NMSE metric, total positron yield is most accurately predicted by the BIC model in Geant4 version 10.2, being (equal) best in 11/15 cases. This is much higher than the next-best combinations (10.1/BIC with 5/11 cases followed by 10/BIC with 4/11). Similar results are observed for 11C yield, with 10.2/BIC achieving (equal) best performance in 14/15 cases, and 10/BIC and 10.1/BIC each achieving (equal) best results in 11/15 cases; QMD also performs reasonably well in this case with 10, 10.1 and 10.2 achieving wins in 6/15 cases. For 10C, 10.6/INCL, 11/INCL and 11.1/INCL are the best performers (each winning in 9/15 cases). Finally, for 15O, 10.2/BIC is the best-performing model with 5/12 wins, followed by 10.4/BIC with 4 wins.
@@
10 9 4 3 3 3 2 4 3 6 3 2 3
10.1 9 4 4 3 3 3 4 3 6 3 2 3
10.2 10 8 4 6 6 3 5 4 5 3 3 3
10.3 9 6 6 6 5 3 2 2 2 4 3 3
10.4 6 8 7 6 5 3 1 1 3 4 3 3
10.5 8 8 6 6 5 4 2 2 3 3 3 4
10.6 9 9 6 6 7 6 2 1 4 4 3 4
10.7 6 6 4 3 5 2 2 1 4 3 2 3
11 8 8 6 4 5 4 3 2 4 3 3 3
11.1 6 5 6 2 3 4 3 2 4 3 3 3
Using the Pearson CCC metric, the best-performing version/model combinations for overall positron yield are 10.2/BIC (10/15 cases), followed by 10/BIC, 10.1/BIC, 10.3/BIC, 10.6/BIC and 10.6/QMD (9/15 cases). Generally, BIC performed very well, with all Geant4 versions achieving (equal) best performance in at least 6 cases. 11C yield was best predicted by 10.6/QMD (7/15 cases) however many version/model combinations did well here also, with 10.2/BIC, 10.2/QMD, 10.3/BIC, 10.4/BIC, 10.5/BIC, 10.6/BIC and 10.6/INCL all achieving 6/15 wins. 10C yield was best predicted by 10/INCL and 10.1 INCL (6/15 cases), closely followed by 10.2/BIC and 10.2/INCL which won in 5/15 cases. The best-performing version/model combinations for 15O yield were 10.3/BIC, 10.4/BIC, 10.5/INCL, 10.6/BIC and 10.6/INCL with 4/15 wins each, and all other version/model combinations achieving 2 or 3 wins.
@@
$\mu$ (mm) $\sigma$ (mm) max (mm) $\mu$ (mm) $\sigma$ (mm) max (mm) $\mu$ (mm) $\sigma$ (mm) max (mm)
10 1 1.85 6 -0.20 1.69 3 1.10 3.82 10.50
10.1 1 1.85 6 -0.20 1.69 3 1 3.91 10.50
10.2 0.60 1.37 3 -0.60 1.58 -3 0.60 3.39 9
10.3 1.30 1.78 6 -0.30 1.41 -3 2.30 4.12 9
10.4 1.60 1.55 6 -0.10 1.44 -3 4 4.45 10.50
10.5 0.69 0.99 3 0.60 2.03 6 2.20 4.24 9
10.6 0.60 1.37 3 -0.10 1.55 3 0.10 2.50 7.50
10.7 1.20 1.72 4.50 0.60 1.95 4.50 0.70 3.40 10.50
11 1.10 1.65 4.50 0.60 1.95 4.50 0.60 3.09 9
11.1 0.30 1.62 3 -0.30 1.62 3 0 2.78 7.50
@@
$\mu$ (mm) $\sigma$ (mm) max (mm) $\mu$ (mm) $\sigma$ (mm) max (mm) $\mu$ (mm) $\sigma$ (mm) max (mm)
10 0.70 1.49 3 0.30 1.41 3 -0.20 1.49 -3
10.1 0.70 1.49 3 0.30 1.41 3 -0.20 1.49 -3
10.2 0.30 1.01 1.50 0 1.13 1.50 -0.50 1.22 -3
10.3 0.50 1.09 3 0.20 1.11 1.50 -0.20 1.37 -3
10.4 0.60 1.11 3 0.30 1.01 1.50 0.10 1.20 -3
10.5 0.35 0.90 1.50 0.20 1.11 1.50 -0.50 1.22 -3
10.6 0.40 0.89 1.50 0.20 1.11 1.50 -0.40 1.33 -3
10.7 1 1.46 3 0.70 1.59 3 -0.10 1.65 3
11 1 1.46 3 0.70 1.59 3 0 1.60 3
11.1 0 1.60 -3 -0.10 1.44 -3 -0.60 1.68 -3
@@ 3.3 Tail region
10 3 2 2 5 9 4 2 2 1 4 4 4
10.1 4 2 2 6 9 4 1 1 2 4 4 4
10.2 12 10 2 11 12 4 1 1 3 7 7 4
10.3 1 1 1 1 1 1 2 4 0 3 4 4
10.4 1 1 1 1 1 1 2 3 0 3 4 3
10.5 1 1 1 1 1 1 3 4 2 2 2 2
10.6 1 2 1 1 1 1 5 3 6 4 10 4
10.7 1 1 1 1 1 1 3 3 5 3 4 3
11 1 1 1 1 1 1 4 3 6 4 4 3
11.1 1 1 1 1 1 1 4 3 5 4 4 3
Using the NMSE metric, 10.2/BIC was the best-performing version/model combination for overall positron yield (12/15 cases), with 10.2/QMD being the second-best (10/15). Results were similar for 11C yield, with the best version/model combinations being 10.2/QMD (12/15 cases) and 10.2/BIC (11/15). For 10C, the most wins were obtained by 10.6/INCL and 11/INCL (6/15 cases) followed by 10.6/BIC, 10.7/INCL and 11.1/INCL (5/15 cases). Finally, for 15O, the best results were obtained with 10.6/QMD (10/12 cases) followed by 10.2/BIC and 10.2/QMD (7/12 cases).
@@
10 12 11 11 10 11 11 4 3 4 8 7 9
10.1 12 11 11 11 11 11 4 4 4 8 7 9
10.2 12 11 11 10 11 11 4 4 4 8 7 7
10.3 11 11 10 11 11 11 2 2 1 8 7 8
10.4 10 10 10 9 11 10 2 2 0 7 7 8
10.5 11 11 11 11 11 11 2 2 3 7 7 8
10.6 11 11 12 11 11 11 2 3 3 7 7 9
10.7 12 11 12 11 11 12 3 3 3 7 7 9
11 12 11 12 11 11 13 3 3 3 8 8 9
11.1 12 11 12 11 11 13 3 3 3 8 8 9
The Pearson CCC results in the tail region were all very similar across Geant4 versions, with only a few wins separating the best and worst-performing version/model combinations in most instances. All version/models exceeded the threshold of 0.95 for a clear majority of cases for total positron yield as well as 11C and 15O production. For total positron annihilation yield, 10/BIC, 10.1/BIC, 10.2/BIC, 10.6/INCL, 10.7/BIC, 10.7/INCL, 11/BIC, 11/INCL, 11.1/BIC and 11.1/INCL all exceeded the target threshold for 12/15 cases. Even the worst-performing version/model combinations still exceeded the threshold in 10/15 cases. For 11C yield, 11/INCL and 11.1/INCL reached the threshold in 13/15 cases (with the worst-performing combination scoring 9/15 wins). Fewer wins were seen with 10C; the best results were obtained with 10/BIC, 10/INCL, 10.1/BIC, 10.1/QMD, 10.1/INCL, 10.2/BIC, 10.2/QMD and 10.2/INCL (4/15 cases). Finally, 15O yield was best predicted by 10/INCL, 10.1/INCL, 10.6/INCL, 10.7/INCL, 11/INCL and 11.1/INCL (9/12 cases) - again, in this case, even the worst-performing version/model combinations exceeded the threshold in 7/12 cases.